OpenStudy (anonymous):

Choose the equation that could be used to find three consecutive integers whose sum is 36. n + (n + 2) + (n + 4) = 36 n + (n + 1) + (n + 3) = 36 n + (n + 1) + (n + 2) = 36 n + (n − 1) + (n − 3) = 36

OpenStudy (anonymous):

Hint: If your first integer is "n", what is the next integer? Isn't that "one more than n"?

OpenStudy (anonymous):

And the integer after the one after "n" is "one more than one more than n". With those 2 hints, you should be able to find the representations of the 3 consecutive integers.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

i just start doin math this year algebra

OpenStudy (anonymous):

Forget about when you started, concentrate on the hints that are given.

OpenStudy (zzr0ck3r):

example of 3 consecutive numbers 1,2,3 if we add them up 1+2+3

OpenStudy (zzr0ck3r):

think about the relation between 2 and 1 which is the same relation between 3 and 2 what is that relation?

OpenStudy (anonymous):

what tha mean to consecutive integers

OpenStudy (anonymous):

Consecutive means the very next one.

OpenStudy (zzr0ck3r):

integers are -3,-2,-1,0,1,2,3..... consecutive means the next one 1,2,3 are 3 consecutive numbers 4,5,6 are 3 consecutive numbers 100212,100213,100214 are 3 consecutive numbers

OpenStudy (anonymous):

For example, the 3 consecutive integers starting with 4 are: 4, 5, 6 That's 4, 4 + 1, 4 + 2

OpenStudy (zzr0ck3r):

^^^^^understand that and you will get it

OpenStudy (anonymous):

So, if you are dealing with "n" instead of "4", just substitute "n" for "4".

OpenStudy (zzr0ck3r):

4,5,6 = 4,4+1,4+2 6,7,8 = 6,6+1,6+2 200,201,202 = 200,200+1, 200+2